OpenStudy (anonymous):

sin(x)^2+4sin(x)-1=0

OpenStudy (anonymous):

what are you trying to calculate?

OpenStudy (anonymous):

4 = cot(x)cos(x)

OpenStudy (anonymous):

we have to fiind the sin of x and then the related angles using a calculator

OpenStudy (anonymous):

Usually we have to factor or use quadratic formula to find sin(x)

OpenStudy (raden):

let t=sin(x), it can be t^2+4t-1=0, can't be factor it so, use quadratic formula :